OpenStudy (anonymous):

i forgot how to solve this type of equation.. 5/9 = t - 1/9. do i add 5/9 to 1/9 which would = 6/9???

OpenStudy (anonymous):

you add 1/9 to both sides of the equation, so it ends up as 6/9 = t (or 2/3 = t simplified)

OpenStudy (anonymous):

okay thanks!